A fighter jet has to hit a small target by flying a horizontal distance when the target is sighted the pilot measures the angle of depression to 30 degree if after 100km the target has an angle of depression of 45degree how far is the target from the fighter jet at that in. stant.

Grade:11

1 Answers

Sudhir Kumar
5 years ago
let horizontal distance be x ang height be h,
tan30 = h/x+100
and tan45 = h/x => x = h 
Put in above equation , 1/31/2 = h/h+100 => h = 100/(1.732 – 1)